OpenStudy (anonymous):

solve the following systems of equations x+3y+4z=14 2x-3y+2z=10 3x-y+z=9

OpenStudy (phi):

As a first step, try adding the first two equations together: x+3y+4z=14 2x-3y+2z=10 can you do that ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3x+6z=24

OpenStudy (phi):

that looks good. Notice all the terms can be divided by 3, and is always nice to simplify you get x + 2z = 8 that is one equation with x and z as unknowns. now we need to use the last equation with one of the first 2 to get rid of the y and give us a second equation. using the first, and the last equation x+3y+4z=14 3x-y+z=9 to get the 3y in the first equation to "cancel out" we need -3y in the last equation to get a -3y in the last equation, multiply both sides (and all terms) by 3. we get 9x -3y +3z= 27 now add the two equations 3x-y+z=9 9x -3y +3z= 27

OpenStudy (phi):

oops, that last line should read now add the two equations x+3y+4z=14 9x -3y +3z= 27

OpenStudy (anonymous):

10x+7z=41

OpenStudy (phi):

now you have the two equations x + 2z = 8 10x+7z=41 you could multiply the first equation by -10, and then add the two equations

OpenStudy (anonymous):

z=-3

OpenStudy (phi):

now pick one of the two equations with x and z (the simpler one is the easiest choice and solve for x)

OpenStudy (phi):

except check your sign on z

OpenStudy (phi):

I got -13 z = -39 z= -39/-13 = +3